### v1

# Perceptions of Linguistic Uncertainty by Language Models and Humans

Abstract
Uncertainty expressions such as “probably” or “highly unlikely” are pervasive in human language. While prior work has established that there is population-level agreement in terms of how humans interpret these expressions, there has been little inquiry into the abilities of language models to interpret such expressions. In this paper, we investigate how language models map linguistic expressions of uncertainty to numerical responses. Our approach assesses whether language models can employ theory of mind in this setting: understanding the uncertainty of another agent about a particular statement, independently of the model’s own certainty about that statement. We evaluate both humans and 10 popular language models on a task created to assess these abilities. Unexpectedly, we find that 8 out of 10 models are able to map uncertainty expressions to probabilistic responses in a human-like manner. However, we observe systematically different behavior depending on whether a statement is actually true or false. This sensitivity indicates that language models are substantially more susceptible to bias based on their prior knowledge (as compared to humans). These findings raise important questions and have broad implications for human-AI alignment and AI-AI communication.
Keywords: Uncertainty Estimation, Large Language Models, Uncertainty Expression, Uncertainty Quantification, Theory of Mind, Human studies
1 Introduction
Uncertainty is ubiquitous in human communication — in relaying predictions (“it is likely to rain tomorrow”), conveying imperfect knowledge (“I think I have a copy in my desk”), and describing unknown information (“the artifact could be more than 500 years old”). Expressing uncertainty is critical in fields such as medicine, law, and politics, where statements including uncertainty expressions (e.g., “likely,” “doubtful”) are frequently used to support medical, judicial, and political decisions (Karelitz and Budescu, 2004). Domain experts use these expressions to communicate uncertainty across a variety of situations, such as the likelihood of side-effects of a medical treatment (Sawant and Sansgiry, 2018; Patt and Dessai, 2005), the chances of winning a not-guilty verdict in legal cases (Fore, 2019), the probability of environmental events resulting from climate change (Patt and Dessai, 2005; Ho et al., 2015), or the likelihood of emergence of military conflicts (Duke, 2023). Prior work has found that, in general, humans are well-attuned to the use of such uncertainty expressions, exhibiting population-level agreement in mapping these expressions to corresponding probabilities (Wallsten et al., 1986a; Willems et al., 2019; Fagen-Ulmschneider, 2019).
However, the topic of how large language models (LLMs) interpret linguistic uncertainty has received relatively little attention (Sileo and Moens, 2023). In particular, given text where a speaker expresses uncertainty about a particular statement, this paper investigates whether LLMs can interpret the uncertainty not as a function of the model internal beliefs, but by objectively assessing the speaker’s uncertainty about the statement. Consider the motivating example in Figure 1: when writing a headline for a statement qualified by the word “probable,” ChatGPT expresses substantially different forms of linguistic uncertainty depending on its prior belief about the statement. [^1] When prompted about its belief about the different statements, ChatGPT agrees with the first one and disagrees with the second; see Figure 9 in the Appendix A. In this example, ChatGPT is conflating the speaker’s uncertainty with its own uncertainty about the statement—in effect, a failure of “theory of mind.”
Figure 1: Two interactions with ChatGPT (June 2024). In each, ChatGPT is asked to write a headline for a short passage. Both passages are structured identically and qualified with the word “probable”, but the first is about climate change and the second about the link between vaccines and autism. For the first passage, ChatGPT generates a confident-sounding headline, using words like “conclude” and “comprehensive”. The second headline is weaker, with words like “suggests” and “possible”.
In this work, we investigate the abilities of LLMs to provide quantitative interpretations of uncertainty expressions, focusing in particular on how the prior knowledge of an LLM affects this ability. To this end, we introduce a new task [^2] All code and data used in this study are available at https://github.com/UCIDataLab/llm-uncertainty-perceptions. in which LLMs must map text containing uncertainty expressions to numerical responses. We analyze the performance of both humans and 10 popular LLMs on this task, enabling direct comparison between humans and models. We find that larger, newer models like GPT-4 and Llama-3 can consistently map uncertainty expressions to numerical responses that align with human population-level perceptions. However, we also show that the responses LLMs generate are susceptible to bias based on their prior knowledge—to a much greater extent than those of humans.
The models’ sensitivity to its prior knowledge has concerning implications given the increasing use of LLMs for generating content (e.g., summarization, data augmentation) (Wang et al., 2023; Shao et al., 2024; Laban et al., 2024), evaluating language generation (Zheng et al., 2023), simulating human behavior (Argyle et al., 2023; Park et al., 2023a), or even in building multi-agent pipelines (Wu et al., 2023; Qian et al., 2024). When an LLM’s ability to quantify uncertainty can be “poisoned” by its beliefs, its downstream performance is dependent on its parametric or pretraining knowledge (which can be obsolete or wrong Liang et al. (2022); Longpre et al. (2023)), rather than on critical contextual information (Longpre et al., 2021). Further, this means that the biases of a model (including the many well-documented potentially harmful biases of LLMs, e.g., Wan et al. (2023); Kotek et al. (2023); Salewski et al. (2024); Scherrer et al. (2024); Motoki et al. (2024)) can subtly manifest in how it interprets and generates uncertainty language and, as a consequence, have broader implications for human-AI and AI-AI interactions.
2 Related Work
Human Perceptions of Uncertainty Expressions.
In fields like medicine, finance, law, and politics, where it is impossible to make predictions with complete certainty, decisions are often informed by subjective probabilities (Karelitz and Budescu, 2004; Dhami and Wallsten, 2005; Fore, 2019). Subjective probabilities can be communicated quantitatively, through numerical probabilities (e.g., odds, percentages, intervals), or qualitatively, through the use of uncertainty expressions or epistemological markers (e.g., “I believe”, “According to”) (Dhami and Mandel, 2022).
Although being less precise than numerical probabilities (Wallsten et al., 1986b; Brun and Teigen, 1988; Budescu et al., 2014), humans generally prefer to use linguistic expressions, rather than numbers, to communicate uncertainty (Erev and Cohen, 1990; Wallsten et al., 1993). Interested in the efficacy of how humans communicate uncertainty linguistically, researchers have examined how individuals map uncertainty expressions into numerical values across different fields and expertise levels (Windschitl and Wells (1996); Karelitz and Budescu (2004); Wallsten et al. (2008, 1986a); Fore (2019); inter alia). Although there can be considerable variation in responses at the individual level, these studies have revealed that there are consistent patterns relating uncertainty expressions and numerical probabilities that can be observed systematically at the population level (Budescu and Wallsten, 1985; Clarke et al., 1992; Wallsten et al., 2008; Willems et al., 2019; Fagen-Ulmschneider, 2019).
Uncertainty Quantification in LLMs.
The need for more reliable LLMs has prompted researchers to investigate new methods for communicating the internal uncertainty of LLMs. Proposed methods differ in the information used to estimate the model’s uncertainty, ranging from methods that use token-level information (Jiang et al., 2021; Kuhn et al., 2023; Duan et al., 2024), to dissimilarities across multiple samples (Si et al., 2022; Chen and Mueller, 2023; Xiong et al., 2024; Hou et al., 2024; Lin et al., 2024; Aichberger et al., 2024), to training external classifiers using the inputs and/or LLMs’ representations (Jiang et al., 2021; Mielke et al., 2022; Shrivastava et al., 2023), or even directly eliciting confidence estimates from LLMs as output tokens (Lin et al., 2022; Tian et al., 2023). Furthermore, several works have analyzed the impact of LLM-articulated uncertainty in human-AI interaction, finding that participants adjust their perception of LLMs’ correctness when shown LLM outputs that include uncertainty expressions (Zhou et al., 2024; Kim et al., 2024; Steyvers et al., 2024). While these works investigate how LLMs express uncertainty when generating text and how humans react to various uncertainty expressions in text, there has been far less work on the questions we focus on in this paper, i.e., how LLMs interpret uncertainty in text and how closely these interpretations match those of humans.
LLM Perceptions of Uncertainty Expressions.
A small body of work has begun to investigate the relationships between uncertainty expressions and model behavior. Sileo and Moens (2023) investigate whether LLMs are able to discriminate between two uncertainty expressions and reason in terms of compositions of expressions. However, the paper focuses on LLMs’ binary rankings of expressions, as opposed to numerical interpretations, and does not compare LLM and human behavior. Most directly related to our work is that of Maloney et al. (2024) which compares numerical probability estimates from GPT-4 and humans using a small set of “context” prompts. Tang et al. (2024) also compare the numeric-textual mapping of uncertainty expressions across four different contexts in both Chinese and English languages. Our paper goes significantly beyond these works by assessing a broad range of LLMs using a more diverse and natural set of contexts. We additionally conduct human experiments, assessing the performance of humans on the same task to facilitate a direct comparison between humans and LLMs. Further, our approach is designed to target “theory of mind”—the task requires humans and LLMs to quantify what an uncertainty expression reflects about the speaker’s belief, rather than what the expression means to them. Finally, our work is the first that we are aware of to investigate how LLMs can be biased by their prior knowledge in mapping uncertainty expressions to numerical probabilities.
3 Baseline Human Study
As a baseline for how people map uncertainty expressions to numerical probabilities, we first conducted an experiment in which 94 humans were shown uncertainty expressions and asked to provide corresponding numerical responses. We focused on a set of 14 uncertainty expressions (e.g., “almost certain”, “unlikely”—the full list is provided in Appendix B and is also shown on the y-axis in Figure 4), drawn from Wallsten et al. (2008) and Wallsten et al. (1986a). In this initial experiment, our goal is to assess how people perceive these uncertainty expressions “in the wild,” putting them in the context of plausible real-world statements. In addition, we also aim to use statements that minimize the potential for people to conflate their own beliefs about these statements with their assessment of the confidence of the person making the statement.
To this end, we constructed a set of statements $(u,s,e)$ which include uncertainty expressions $u\in\mathcal{U}$ used by speakers $s\in\mathcal{S}$ to convey their degree of certainty about the truthfulness or falsehood of a statement or event $e\in\mathcal{E}$. By presenting statements about a specific speaker $s$, [^3] Names are selected arbitrarily from a pre-defined pool of names. Participants see each name once throughout the experiment. For more details, see Appendix B. we are asking participants to use “theory of mind” to estimate how likely is that the speaker believes that the statement is true. We then query participants about the speaker’s strength of belief, clearly distinguishing this notion from the participant’s own beliefs. For instance, given the statement “Sonia believes it is unlikely it will rain today,” we can ask participants to quantify with a numerical response how likely Sonia thinks it is that it will rain today, distinct from the probability the participants themselves believe it will rain. We use the term numerical response throughout the paper to refer to the participant’s numerical assessment of the likelihood that speaker $s$ believes statement $e$ to be true when using uncertainty expression $u$. We instruct participants that responses should be numbers between 0 and 100, where 0 implies that a speaker $s$ believes there is a $0\%$ chance that the statement $e$ is true whilst 100 implies a speaker believes that there is a $100\%$ chance that the statement is true.
In our baseline experiment, we use non-verifiable statements $e$ to separate the meaning of the uncertainty expressions from uncertainty about the statements themselves. These are statements that are not sufficiently grounded with specific contextual information to allow an external observer to be confident in either the truth or falsity of the statement. For example, in the context of a prompt such as “Maria believes it is likely that [statement]”, statements such as her boss has two pets or her flight will land around 6pm are statements we consider non-verifiable in the sense that there is insufficient context provided (that the speaker knows, but does not state) for an observer to be able to reliably assess the likelihood that the statement is true. In contrast, verifiable statements (which we discuss further in Section 4.1) can be verified as correct or incorrect in a context-free sense (e.g., the capital city of Peru is Lima); humans and LLMs will often have strong prior beliefs about the likelihood that such statements are true.
Specifically, we manually constructed a set of 60 non-verifiable statements and systematically combined these with 14 uncertainty expressions. We randomly generated speaker names, generating sentences describing the belief of a hypothetical speaker in the form: “[Speaker] believes it is [uncertainty phrase] that [statement].” For each of these sentences, participants were asked to quantify the speaker’s belief about the statement, in particular, they were asked what is the probability being expressed from the speaker’s perspective that the statement is correct. Participants then provided their response quantized to numerical bins $0,5,10,\ldots,95,100$. Figure 2 illustrates one of the examples shown to participants in the experiment. Each of the 94 participants provided annotations for 28 randomly chosen statements (and speaker names). As a result, every uncertainty expression was annotated twice by each participant.
The outcome of this experiment [^4] Additional details about the experiments can be found in Appendix B. is an empirical distribution over the $2\times 94$ numerical responses that participants associated with each uncertainty expression. For example, Figure 3 reflects the histogram of responses assigned to the phrases “very likely” and “very unlikely”; results for all 14 uncertainty expressions can be summarized in a heat-map as shown in Figure 4. Overall, we observe results that are in agreement with prior work on human perceptions of uncertainty expressions (Wallsten et al., 2008, 1986a; Willems et al., 2019), including consistent ordering in aggregate population patterns, as determined by the mode of the empirical distributions.
Figure 2: Example question of a non-verifiable statement provided to participants in the baseline experiment. Each example constitutes a unique name and statement. Participants see one question at a time.
(a) “very likely”
(b) “very unlikely”
Figure 3: Histogram of participant responses for non-verifiable statements for two uncertainty expressions.
Figure 4: Human empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting. Highlighted blue boxes represent the mode of each expression.
4 Methodology
Our full set of experiments includes experiments with both humans and LLMs as participants and with both non-verifiable (NV) and verifiable (V) statements. The baseline experiment, as described in Section 3, consists of human participants and non-verifiable statements (denoted human + NV). In this section, we describe extensions to include verifiable statements and LLM participants, resulting in three additional sets of experiments: human + V, LLM + V, LLM + NV. To help us draw comparisons between the three additional settings and the baseline experiment, we conclude this section with the description of two quality metrics.
4.1 Verifiable Statements
In addition to the non-verifiable statements described in Section 3, our dataset also includes verifiable statements, for the purpose of assessing the effects of prior knowledge on quantifying linguistic uncertainty. To this end, we create 60 verifiable statements based on a multiple-choice question-answering dataset from The Question Company.555 https://www.thequestionco.com/ Starting with 30 of the dataset’s “easy” questions and corresponding multiple-choice options, we write true statements that use the correct answer and false statements using one of the incorrect answers. Examples of verifiable statements and additional details about the dataset are included in Appendix B. We analyze results for both humans and LLMs with these 60 verifiable statements in the main paper: for additional validation, our dataset includes 400 additional statements (generated from the AI2-ARC question answering dataset (Clark et al., 2018) via a similar procedure), which we analyze in Appendix E.
4.2 Numerical Responses from LLMs
To obtain uncertainty estimates from LLMs, we create prompts similar to the queries provided to humans (see list of prompts in Appendix B.4). Our goal is to estimate an empirical distribution, per uncertainty expression $u$, over the LLM’s generated responses, in a manner similar to how empirical distributions for humans are generated (e.g., see Figure 4). In the results in this paper we focus on greedy decoding, where we select the numerical response that has the highest probability in the next-token probability distribution generated by the LLM conditioned on the prompt, i.e., decoding with temperature=0. Because this sampling approach requires no knowledge about the weights or next-token probabilities, it is applicable to any model, including those behind black-box APIs, such as Gemini (Anil et al., 2024) and GPT-4 (Achiam et al., 2024). We focus primarily on this greedy sampling approach since it aligns more closely with the human responses, i.e., we want to assess the ability of each LLM as if it were an individual human providing a single response rather than asking it to match a population distribution of human responses [^6] In Section 5 we evaluate the sensitivity of our conclusions by also analyzing decoding with temperature=1.. We obtain an empirical distribution of LLM responses, for each uncertainty expression $u$, by generating responses in this manner over multiple statements $e$ and speakers $s$. We refer the interested reader to Appendix C for additional information on the extraction methodologies used.
4.3 Metrics
We treat the empirical distribution obtained for the non-verifiable statements with human participants (described in Section 3) as our reference distribution for evaluation purposes, since it reflects human perceptions of uncertainty expressions in a setting that is designed to be free of prior information or biases about the corresponding statements. For every uncertainty expression $u\in\mathcal{U}$, we define a reference conditional probability distribution $P(k|u)$, $k=0,5,10,\ldots,95,100$. where $P(k|u)$ is the empirical distribution from the baseline experiment. Given a response from any agent, human or LLM, in the context of a particular uncertainty expression $u$ we measure the quality of the response using the reference distribution $P(k|u)$.
The primary quality metric that we propose is Proportional Agreement (PA). PA can be defined as follows: if an agent’s response matches bin $k$ for uncertainty expression $u$, then the PA value for that response is defined as $P(k|u)$, where $P$ is the reference (population) distribution defined above. Intuitively, for an expression $u$, this PA score $P(k|u)$ represents the probability that the agent’s response $k$ agrees with that of a randomly selected individual, and is upper bounded for any expression by $\arg\max_{k}P(k|u)$, i.e., by the mode of the $P(k|u)$. The higher the PA value, the better the quality of the response in terms of agreement with the aggregate human population (as reflected by $P(k|u)$). To compute a single score for a particular LLM or individual human, we average the PA score over multiple responses and over the 14 uncertainty expressions [^7] Note that the PA metric is similar to the log-probability metric widely used to score probabilistic models in machine learning. However, it is not a likelihood in the sense that a likelihood corresponds to measuring the probability mass a model assigns to an observed outcome. Thus, in this non-likelihood context it is appropriate to average the PA scores directly (rather than taking products of probabilities as would be done under an IID likelihood assumption)..
An alternative to the PA metric would be to compare histograms of responses, e.g., based on multiple responses from agents for a particular uncertainty expression $u$. We provide numerical results for histogram comparisons (using the Wasserstein distance between histograms) in Appendix D, but we consider this to be of secondary interest since we are not requiring any LLM or individual human to necessarily replicate the full population variability of responses.
As an additional measure of alignment between the reference distribution and the agent’s distribution, we also compute the Mean Absolute Error (MAE), for each uncertainty expression $u$, defined as the absolute difference between (i) the mean of the responses across statements involving $u$ for an agent, and (ii) the mean of the reference distribution for $u$, $P(k|u)$. We then average across the 14 expressions $u$ to get a single score per agent.
5 Results
This section examines the ability of several well-known LLMs to interpret uncertainty expressions [^8] We focus in this section on a subset of popular models; results for all 10 models evaluated are in Appendix D.. We begin by assessing models’ abilities to produce numerical responses that resemble human-like trends (e.g., higher numerical responses assigned to higher certainty expressions and vice-versa). We then study the effect of prior knowledge in the perception of uncertainty of both humans and models. We conclude with an assessment of the generalizability of our findings.
5.1 How well do LLMs perceive uncertainty?
As established in prior work (Wallsten et al., 1986b; Fagen-Ulmschneider, 2019; Willems et al., 2019) and in our baseline experiment (Section 3), humans show population-level agreement in mapping uncertainty expressions to numerical responses. In this section, we assess whether LLMs possess a similar ability to ascribe numerical responses to uncertainty expressions. To this end, we prompt LLMs to provide responses for the same non-verifiable (NV) statements as in the baseline experiment. Figure 5 shows the expression-wise histograms for these responses for two LLMs: GPT-4o and OLMo (7B), each of which can be directly compared to the histogram for humans in Figure 4.
(a) GPT-4o
(b) OLMo-7B Instruct
Figure 5: Model empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting (LLM + NV). Highlighted boxes represent the mode value for each expression. Even though we found no evidence of explicit instruction tuning datasets focusing on uncertainty estimation tasks, our results show that GPT-4o generally manifests human-like behavior, whereas OLMo (7B) does not.
Visually, we observe that GPT-4o matches the human distributions well, with smaller variance per distribution, while OLMo (7B) is less well-matched. In Figure 13 in Appendix D, we see that most LLMs map uncertainty expressions to numerical responses in a manner consistent with human behavior, with higher values for expressions that are perceived by humans as higher-certainty (e.g., “almost certain,” “highly likely”) and lower values for lower-certainty expressions (e.g., “very unlikely”). Only two of the LLMs evaluated, OLMo (7B) and Gemma (2B), fail to reproduce this “increasing” pattern across expressions (relative to the human pattern). One clear difference between humans and LLMs is that the conditional distributions of LLMs have lower entropy (or variance) relative to the human distributions, with the LLM distributions tending to be much more concentrated on a small number of responses compared to the variance in responses from a population of humans.
These observations are reflected more precisely by the PA scores in Table 1. We observe that larger and newer LLMs (in particular, GPT-4, LLama3 (70B), and Gemini) perform especially well on this task under the PA metric, being at 85% or above in terms of matching the modal scores that a human population assigns to each uncertainty expression. In fact, 7 out of the 10 LLMs evaluated are significantly better matched to population modal responses than are individual humans on average [^9] The average performance of individual humans is represented by the Human Individual row in Table 1.. This aligns with the high-level findings of Maloney et al. (2024), in particular, that the difference between the numerical responses of GPT-4 and humans were similar to (or smaller than) inter-human differences. In the context of our experiments, these high scores reflect that LLMs tend to be more consistent than individual humans in terms of agreement with aggregate human responses.
The MAE scores in Table 1, where lower is better, are highly anti-correlated with the PA scores, and tell a similar story in terms of which models perform better. To provide a sense of scale, the MAE numbers are lower-bounded by 0 and upper-bounded by 25 (the expected MAE for random responses).
Table 1: Human-LLM agreement for non-verifiable statements. Average Proportional Agreement (PA), PA as a fraction of the Human Mode results (PA), and absolute error between mean responses (MAE). Human Mode represents the mode of the human NV distribution, whereas Human Individual represents the PA score of individual human responses relative to the population. 95% confidence intervals were computed using the adjusted bootstrap percentile method.
PA (95% CI) % PA MAE (95% CI)
Human Mode 27.6
Human Individual 17.6 (17.2, 18.0) 63.8 8.91 (8.08, 9.17)
[…]
GPT-4 24.4 (24.2, 24.6) 88.4 4.64 (4.53, 4.74)
GPT-4o 18.9 (18.4, 19.5) 68.5 5.58 (5.34, 5.79)
Gemini 25.4 (25.1, 25.6) 92.0 4.09 (3.92, 4.23)
Llama3 (8B) 17.8 (17.4, 18.2) 64.5 11.99 (10.92, 13.13)
Llama3 (70B) 23.6 (23.4, 23.8) 85.5 5.56 (5.34, 5.80)
Mixtral 8x7B 21.8 (21.4, 22.0) 79.0 5.88 (5.44, 6.34)
Mixtral 8x22B 21.8 (21.6, 22.0) 79.0 7.20 (6.56, 7.81)
OLMo (7B) 11.1 (10.7, 11.5) 40.2 21.41 (20.17, 22.41)
Gemma (2B) 8.1 (7.7, 8.5) 29.3 20.17 (19.11, 21.17)
5.2 Does knowledge affect uncertainty perceptions of LLMs?
In this section, we assess the extent to which LLMs, and humans, are biased by their prior knowledge or beliefs in mapping uncertainty expressions to numerical responses. To investigate this question we collect responses from humans and LLMs on our verifiable (V) dataset, which includes both true and false common-knowledge statements (based on correct or incorrect answers, respectively, to multiple-choice questions). We find that average PA scores for both humans and LLMs are systematically lower for verifiable statements compared to the non-verifiable responses (Table 2). This suggests that prior knowledge about a statement makes it more difficult to quantify the beliefs of someone else about that statement. While humans show a small drop in their PA score, this reduction in PA is particularly pronounced for LLMs: all 10 LLMs demonstrated a significant reduction in PA, averaging a 4.3 point drop in score (across all models), compared to a 0.9 point drop for humans.
Table 2: Human-LLM agreement for verifiable statements. Average Proportional Agreement (PA), absolute error between mean responses (MAE), and the difference between these scores and those from the non-verifiable statements (Table 1) ($\Delta$ PA and $\Delta$ MAE, respectively). Again Human Mode represents the mode of the human NV distribution, whereas Human Individual represents the average behavior across individual humans on the verifiable setting.
PA (95% CI) $\Delta$ PA MAE (95% CI) $\Delta$ MAE
Human Mode 27.6
Human Individual 16.7 (16.3, 17.1) -0.9 9.35 (8.23, 9.50) 0.44
[…]
GPT-4 22.1 (21.7, 22.4) -2.3 3.84 (3.03, 4.45) -0.80
GPT-4o 15.2 (14.5, 15.9) -3.7 7.05 (6.45, 7.62) 1.47
Gemini 21.3 (20.7, 21.8) -4.1 7.23 (6.04, 8.49) 3.14
Llama3 (8B) 10.1 (9.5, 10.7) -7.7 16.59 (15.10, 18.18) 4.60
Llama3 (70B) 18.9 (18.1, 19.6) -4.7 13.73 (11.92, 15.64) 8.17
Mixtral 8x7B 15.2 (14.5, 15.9) -6.6 12.23 (10.37, 14.18) 6.35
Mixtral 8x22B 18.6 (18.3, 19.0) -3.2 9.78 (8.51, 11.08) 2.58
OLMo (7B) 7.6 (7.2, 8.0) -3.5 33.66 (31.82, 35.11) 12.25
Gemma (2B) 5.3 (5.0, 5.6) -2.8 25.07 (23.45, 26.66) 4.9
To investigate these differences in more detail, we consider the mean response values produced by the 6 models exhibiting the highest PA score. These values differ systematically depending on whether the statement is true or false: across the 6 LLMs in Figure 6, the mean response is 7.0 percentage points lower for false than true statements. This indicates that the models assign higher response values to the same uncertainty expression when they believe the associated statement it refers to is true than when they believe it to be false (providing a quantitative validation of the ChatGPT example in Figure 1 in Section 1).
Figure 6: Mean numerical response for the verifiable statements discriminated by truthfulness of statements. The mean numerical responses produced by LLMs when evaluated in the context of true statements is significantly larger than when evaluated in the context of false statements. This difference is much larger in magnitude than the difference shown by a human population.
Results for a subset of specific uncertainty expressions are shown in Figure 7. We observe that the prior-knowledge bias is remarkably different depending on the uncertainty expression: the difference between true and false statements is much higher (49.5 percentage points) for the expression “possible” than for the expression “uncertain” (for this expression most models are close to the average 5.7 percentage point difference).
(a) “possible”
(b) “uncertain”
(c) “unlikely”
Figure 7: Mean response for verifiable statements discriminated by truthfulness of statements for three different uncertainty expressions. We observe that the differences in mean numerical responses differs by uncertainty expression. Despite changes in magnitude, ChatGPT systematically exhibits the largest gap.
Overall, we find that all of the evaluated LLMs demonstrate significant biases based on their prior knowledge, well beyond those of humans. Our results indicate that when an LLM believes a statement is false, it tends to perceive the speaker’s certainty as low, regardless of the actual uncertainty expressed by the speaker (and vice versa), lacking “theory of mind." Our findings in general are also consistent with recent results in the context of LLM reading comprehension, where the efficacy of LLM comprehension is sensitive to the degree of prior knowledge it has about the text it is analyzing Basmov et al. (2024).
5.3 How generalizable are our findings?
In the previous sections, our analyses are conducted on a manually curated set of 120 statements, comprised of 60 NV statements and 60 V statements. To further validate our findings concerning LLMs’ prior knowledge biases, we re-assess the impact of knowledge in LLMs’ perceptual capabilities by obtaining their responses for 400 additional verifiable statements. Similarly to the original study, Figure 15 shows that, on average, all models except Gemma (2B) exhibit significant perceptual differences between true and false statements— between 5.87 (OLMo (7B)) and 17.26 (LLama3 (70B)) percentage points. While the magnitude of the difference is different across the two datasets (potentially due to semantic differences between the two QA datasets used to curate the verifiable statements), the directionality of the results with this larger dataset nonetheless corroborates our knowledge bias finding by showing that these perceptual differences persist in a different context. We refer the reader to Appendix E for a more detailed description of the experimental setup and additional results.
5.4 How does decoding impact our findings?
The previous analyses employ greedy decoding (i.e., temperature=0) when obtaining numerical responses from the LLMs. In this section, we investigate the impact of the decoding technique in the model’s abilities to perceive linguistic uncertainty, by considering richer probability information (i.e., temperature=1) when obtaining the response [^10] This analysis requires full next-token probability information from an LLM, which is prohibitively expensive to obtain empirically through sampling as it would require a large sample size (per $(u,s,e)$) to obtain accurately. As a result, we limit our analysis to OpenAI models for which the top 20 next-token probabilities are available. See Appendix F for additional discussion on this topic..
Table 3 summarizes the change in agreement between LLM and human responses between the verifiable and non-verifiable settings (in terms of change in PA and MAE) when using probabilistic decoding. Validating the results reported in Section 5.2 with greedy decoding, we observe a clear difference in the PA score between non-verifiable and verifiable statements when using probabilistic decoding. Further, comparing responses across true and false statements, we observe large mean response differences of 11.4, 11.7, and 27.9 percentage points for GPT-4o, GPT-4, and ChatGPT, respectively (see Figure 8 and Figure 17 in the Appendix for a breakdown across expressions). Although GPT-4o mean responses are considerably lower than in the greedy decoding setting (with a 20 percentage points drop), the gap between true and false statements persists. Ultimately, this analysis confirms the robustness of our previous findings to the decoding strategy.
Table 3: Differences in average proportional agreement and mean responses from non-verifiable to verifiable settings when considering probabilistic decoding (temperature=1). Even with a different decoding, we observe the same decrease in LLM perceptions when comparing non-verifiable with verifiable settings.
$\Delta$ PA $\Delta$ MAE
ChatGPT -3.6 0.4
GPT-4 -3.0 1.9
GPT-4o -4.2 -0.6
Figure 8: Mean response on the verifiable statements discriminated by truthfulness of statements when decoding probabilistically temperature=1. Despite being to a less extent, the prior-knowledge bias remains larger than that exhibited by human population.
6 Discussion
Theory of mind.
A growing body of work aims to assess the theory of mind capabilities of LLMs in different contexts (e.g., (Street et al., 2024; Verma et al., 2024; Sap et al., 2022; Zhou et al., 2023)). Our task of mapping uncertainty expressions to numerical probabilities, from the perspective of some speaker, is one component of a general theory of mind ability. Our results indicate that LLMs have room for improvement in this area, in particular, that they are prone to confusing their own belief about a statement with the belief of someone else.
Connection to human behavior simulation using LLMs.
Our experiments reveal that, despite agreeing with population-level perceptions of linguistic uncertainty, models do not capture the full diversity of human behavior. Given the recent interest in using LLMs to simulate human participants (Aher et al., 2023; Gui and Toubia, 2023; Dillion et al., 2023; Park et al., 2023b; Namikoshi et al., 2024), our work raises important questions about whose opinions and behaviors are being simulated (Santurkar et al., 2023; Motoki et al., 2023) and reveals a new dimension in which human and model diversity differ.
Limitations
Biases in Interpreting Uncertainty Expressions: Prior work has raised several concerns about the consistency of humans’ interpretations of uncertainty expressions, demonstrating that they are subject to a number of biases and nuances. For example, people may conflate the speaker’s confidence with the speaker’s estimated uncertainty (Fleiner and Vennekens, 2024), statements worded in terms of confidence (“I am almost certain”) or likelihood (“I believe it is almost certain”) are interpreted as primarily communicating different types of uncertainty (epistemic and aleatoric, respectively) (Ülkümen et al., 2016), and these statements are directional, emphasizing either the occurrence or non-occurrence of an event (Teigen, 2023). Further, these interpretations are context-dependent, affected by factors including the individual’s perception of the speaker and the severity of the event in question (Budescu and Wallsten, 1985; Collins and Hahn, 2018; Bonnefon and Villejoubert, 2006; Juanchich and Sirota, 2013; Brun and Teigen, 1988; Weber and Hilton, 1990). Thus, in interpreting uncertainty expressions, high variability can occur both across (Zhang et al., 2023; Collins and Hahn, 2018) and within (Clarke et al., 1992; Van Der Bles et al., 2019) individuals. In this paper we do not explore these dimensions of the interpretation of uncertainty expressions, which could limit the generality of our conclusions.
US Centric View: In this paper we focus on a small set of uncertainty expressions in English and our baseline is drawn from participants located in the United States. Investigating the role that cultural and language differences play in communicating uncertainty is important future work that will help better characterize the downstream abilities of LLMs for all users. The recent work by Tang et al. (2024) represents a first step in this direction, but it does not account for the cultural context that would inform the meaning of a speaker (Huang and Yang, 2023). In particular, the meaning of “probable” and “very likely” may differ significantly between English and Chinese. Such cultural difference may suffice to explain the observed differences in perceptions measured between US population and GPT-4 when prompted with Chinese contexts.
Lack of Explanation: Our results highlight the LLMs’ abilities to interpret uncertainty phrases in a way that agrees with population-level human distribution in the non-verifiable and to less extent in the verifiable setting. It is not clear why models have acquired this general ability, given that there are not similarly framed tasks in available instruction-tuning and human feedback datasets (Wang et al., 2022; Bai et al., 2022). We hope that future work will explore the origins of this behavior.
7 Conclusions
We introduce the task of assessing the abilities of LLMs to interpret uncertainty in language and evaluate a number of models in this context. We observe that many LLMs can competently map uncertainty expressions to numerical responses in a way that aligns with population-level human perceptions, although the probabilities they choose are much less diverse than those by humans. Additionally, we find that LLMs are more susceptible to conflating their own uncertainty about a statement with the statement speaker’s uncertainty, resulting in performance that is biased by the LLM’s belief about the statement.
In proposing this task, we do not take a stance on whether LLM behavior should mirror the diversity of human behavior—which is a broader philosophical discussion—but focus on characterizing LLMs in comparison to human patterns that arise at the population-level. By highlighting systematic inconsistencies related to the perceptions of linguistic uncertainty in the presence of knowledge, we shed light into overlooked model behaviors that are critical for understanding human-AI communication and downstream LLM performance.
Appendix A Motivating example
In the main paper, we discuss the divergence of model behavior in the use case of news headline generation when prompted with text containing uncertainty expressions. In particular, Figure 1 shows two different levels of conviction in the generated text: the first headline contains confident-sounding words like “conclude” and “comprehensive” whereas the second generated headline uses less confident language like “suggests” and “possible”. To further understand the extent to which these differences could be explained by the models’ own knowledge, we prompt the model a second time to assess its agreement or disagreement with the topic being discussed. The answers are illustrated in Figure 9.
Figure 9: Two interactions with ChatGPT (June 2024). When asked about its belief in the statements from Figure 1. ChatGPT agrees that human activities are the main driver of climate change, but disagrees with the statement that early vaccination increases the chances of developing autism.
Appendix B Experiment Details
This section describes in greater detail various aspects of the experiments conducted in this paper, including the list of uncertainty expressions, the name selection strategy, the list of prompts, a list of statements, as well as additional details on the human experiments.
B.1 Uncertainty Expressions
The uncertainty expressions are a subset of the expressions proposed in Wallsten et al.; Wallsten et al.; Willems et al.; Fore. The final list of uncertainty expressions used in this paper is listed below:
  • almost certain, highly likely, very likely, likely, probable, somewhat likely, somewhat unlikely, uncertain, possible, unlikely, not likely, doubtful, very unlikely, highly unlikely
B.2 Name Selection
All names used in our experiments were collected from a random name generator [^11] https://randomwordgenerator.com/name.php, last accessed on March 26th, 2024., which we ran iteratively until we obtained 32 unique names, half of each biological gender (as determined by the random generator).
  • Female names: “Amanda”, “Bonnie”, “Camille”, “Catherine”, “Cheri”, “Ethel”, “Gabriela”, “Jacquelyn”, “Jessica”, “Laura”, “Olga”, “Roxanne”, “Silvia”, “Tara”, “Violet”
  • Male names: “Brendan”, “Bruce”, “David”, “Gary”, “Isaac”, “Jeffery”, “Joey”, “Johnnie”, “Kenny”, “Lance”, “Marco”, “Mike”, “Nathan”, “Nick”, “Raul”
B.3 Human Experiments
Human responses were collected using Prolific (https://www.prolific.com/). We recruited 100 participants for the non-verifiable experiment and 100 different participants for the second verifiable experiment. One of the 100 responses was not received due to a technical issue in both the first and second experiment, leaving a total of 99 responses for each. We recruited participants whose first language was English that were located in the United States. Participants were paid \$2 for completing the study and the average completion time was 8 minutes and 48 sections; the average payment rate was \$13.64/hour. The University of California, Irvine Institutional Review Board (IRB) approved the experimental protocol. Prior to the experiment, participants were given detailed instructions outlining the experimental procedure as well as how to understand and interact with the user interface. Participants were asked to sign an integrity pledge after reading all of the instructions, stating that they would complete the experiment to the best of their abilities. After submitting their integrity pledge, participants were granted access to the experiment.
We filtered out low-quality responses with the following procedure. For each participant, we computed the Spearman correlation between the participant’s responses and the overall ranking of uncertainty statements in the non-verifiable experiment. We removed participants with $\rho<0.2$, a threshold chosen empirically to filter out only no-signal, spam-like responses. This filter removed 5 participants in the first experiment and 10 in the second experiment. In total, we remain with 94 participants in the non-verifiable experiment and 89 in the verifiable experiment.
B.4 Prompts
In our main paper, we conduct experiments using 2 demonstrations, since this better replicates the human setup.
In this experiment, you will be shown text reflecting an individual’s beliefs about a statement. You will then be asked to judge, in your opinion, the strength of that individual’s belief. To do so, you will provide a probability (expressed as a percentage) where: - The number 0 indicates that the individual believes with certainty that the statement is false. - The number 100 indicates that the individual believes with certainty that the statement is true. You will be shown two example question and response pairs below to familiarize you with the experiment setup. After the examples, you will be asked to answer a new question. ## Example Question 1 [[demonstration1]] Choose the answer from the following options: 0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer: 100 ## Example Question 2 [[demonstration2]]|
Choose the answer from the following options:
0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100
Correct answer: 100
## Question
Given the examples before, answer the following question by writing a single number as the answer.
[[test-example]]
Choose the answer from the following options:
0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer:
Figure 10: Prompt used to conduct the experiments in the 2-shot setting. [[demonstration1]], [[demonstration2]], and [[text-example]] are placeholders that will be composed of a statement (or event), question, and individual instructions. The prompt closely reflects the interface presented to the participants of the user studies. Similarly to the human interface, we encourage LLMs to select a multiple of 5 in the range [0, 100].
Sentence: “[[speaker]] believes it’s [[uncertainty]] that [[statement]].” Question: Provide a number between 0 and 100 that reflects, from [[speaker]]’s perspective, the probability (expressed as a percentage) that [[statement]]. - 0 indicates [[speaker]] believes there is a 0% chance that [[statement]]. - 100 indicates [[speaker]] believes there is a 100% chance that [[statement]].
Figure 11: Structure of a demonstration (or exemplar). The placeholders [[speaker]] are replaced by gendered names, [[uncertainty]] is replaced by the uncertainty expressions, and [[statement]] is replaced with the corresponding statements.
During the course of our experiments, we carried experiments with varying assumptions: non-verifiable setup assessed models (and humans) perceptions in the absence of strong prior knowledge about the statements, whereas the verifiable setup focused on the evaluation of the same perceptions when knowledge was present. We used two different sets of exemplars in our experiments to reflect these differences. The set of exemplars used in the non-verifiable setting are defined in terms of the following:
  • speaker: “Kathleen”, uncertainty: “impossible”, statement: “the cafe made a profit in the last 6 months”.
  • speaker: “Cedric”, uncertainty: “certain”, statement: “the new treatment will improve the patient’s condition”.
As for the verifiable exemplars, we experimented with various orderings of the statements and found minimal differences in the obtained distribution of numerical responses. As a result, we report the results using a false statement as the first example and a true statement as the second example in the prompt. These were associated with the following speaker names and uncertainty expressions:
  • speaker: “Kathleen”, uncertainty: “impossible”, statement: “the Sun orbits around the planet Earth”.
  • speaker: “Cedric”, uncertainty: “certain”, statement: “all metals can conduct electricity”.
B.5 Non-Verifiable Statements
In this section, we list a random selection of 20 statements (5 per each statement type). These statements are incorporated in the prompts listed in the main paper and the placeholders [[they]] and [[their]] are replaced by pronouns matching the gender of the statement speaker’s name.
Forecasting of future events.
Verbal probabilities are often used to communicate uncertainty about future events.
  1. 1.
    [[they]] will buy a new watch this Thanksgiving weekend.
  2. 2.
    [[they]] will be offered a promotion this fall.
  3. 3.
    the company will have another round of lay-offs by mid July.
  4. 4.
    there will be vegetarian options at the barbecue.
  5. 5.
    [[they]] will visit New York over winter break.
Imperfect knowledge.
Verbal probabilities can also be used to communicate uncertainty imprecise information about events or outcomes.
  1. 1.
    the restaurant near [[their]] apartment accepts reservations.
  2. 2.
    the new museum is offering complimentary admission.
  3. 3.
    there is a yoga studio within 2 miles of [[their]] workplace.
  4. 4.
    there are more than eighty students in the auditorium right now.
  5. 5.
    the temperature in the office is at least 72 degrees Fahrenheit.
Possession.
Alternatively, verbal probabilities can be used to convey uncertainty about acquaintances, be it in terms of the objects they own or in terms of their preferences.
  1. 1.
    [[their]] boss owns a blue car.
  2. 2.
    [[their]] friend has a leather jacket.
  3. 3.
    [[their]] cousin has a vegetable garden.
  4. 4.
    [[their]] classmate owns a guitar.
  5. 5.
    [[their]] boss has a stereo amplifier.
Preference.
  1. 1.
    [[their]] cousin prefers spinach over broccoli.
  2. 2.
    [[their]] boss prefers coffee over tea.
  3. 3.
    [[their]] friend prefers running over cycling.
  4. 4.
    [[their]] neighbor prefers the beach over the mountains.
  5. 5.
    [[their]] coworker prefers reading books over watching movies.
B.6 Verifiable Statements
In this section, we list a random selection of 9 true statements (3 per each topic) and their false counterparts. These statements are incorporated then used as part of the test examples the prompts listed in Section B.4.
Geography.
One of the topics of the experiment involves geography, as well as knowledge about landmarks and monuments. These statements were curated from a set of easy trivia questions provided by The Question Company (as described in Section 4). For each question-answer pair in the trivia dataset, we create both a true and a false statement using the correct and one incorrect answer choice, respectively.
Given our interest in attesting the knowledge effect in the models’ and humans’ perceptions of linguistic uncertainty, we purposely decided to use easy trivia questions as the basis for our facts (as opposed to more difficult ones), since this subset constitutes a good proxy for facts that LLMs and humans may have strong prior beliefs about.
  1. 1.
    Great Britain directly borders 0 countries.
  2. 2.
    the Colosseum, a famous landmark in Rome, was originally built as an Amphitheatre.
  3. 3.
    New York is known as the Big Apple.
  4. 4.
    Great Britain directly borders 2 countries.
  5. 5.
    the Colosseum, a famous landmark in Rome, was originally built as an Cathedral.
  6. 6.
    New York is known as the Big Orange.
History of Art.
One of the topics of the experiment involves history of arts. For each fact we include both a true and one false variation of that fact.
  1. 1.
    the Mona Lisa is a famous painting by Leonardo da Vinci.
  2. 2.
    the Scream is the best known painting by Edvard Munch.
  3. 3.
    Andy Warhol became a famous artist in the 1960s for painting soup cans and soap boxes.
  4. 4.
    the Mona Lisa is a famous painting by Tintoretto.
  5. 5.
    the Scream is the best known painting by Jackson Pollock.
  6. 6.
    Frida Kahlo became a famous artist in the 1960s for painting soup cans and soap boxes.
Science.
These include facts concerning chemistry, biology, and astronomy. For each fact we include both a true and one false variation of that fact.
  1. 1.
    water’s chemical formula is H2O.
  2. 2.
    pH is a measure of the acidity or basicity of a substance.
  3. 3.
    the nearest planet to the sun is Mercury.
  4. 4.
    carbon monoxide’s chemical formula is H2O.
  5. 5.
    oG is a measure of the acidity or basicity of a substance.
  6. 6.
    the nearest planet to the sun is Mars.
B.7 Language Models
Throughout our paper, we use OpenAI to obtain the results for ChatGPT, GPT-4, and GPT-4o; Google’s Vertex AI APIs to obtain results for Gemini, TogetherAI [^12] https://www.together.ai/ to run LLama3 (70B), Mixtral 8x7B, and Mixtral 8x22B. We run LLama3 (8B) OLMo (7B) and Gemma (2B) locally on a single GPU 8 RTX A6000 (48 GB).
During the paper, we shorten the name of the studied models for simplicity. All our experiments consider the instruction-tuned or RLHF version of the mentioned models. All experiments were conducted from April through June. For reproducibility, we list below the mapping from model name to exact version of the model used:
  • ChatGPT: gpt-3.5-turbo-0125
  • GPT-4: gpt-4-turbo-2024-04-09
  • GPT-4o: gpt-4o-2024-05-13
  • LLama3 (8B): meta-llama/Meta-Llama-3-8B-Instruct
  • LLama3 (70B): meta-llama/Llama-3-70b-chat-hf
  • Gemini: models/gemini-pro
  • Mixtral 8x7B: mistralai/Mixtral-8x7B-Instruct-v0.1
  • Mixtral 8x22B: mistralai/Mixtral-8x22B-Instruct-v0.1
  • Gemma (2B): google/gemma-1.1-2b-it (we found Gemma (2B) to respond better empirically to the prompts than its 7B version, which tended to extrapolate the few-shot instructions with additional examples).
  • OLMo (7B): allenai/OLMo-7B-Instruct
Appendix C Extracting LLMs Numerical Responses
In this section, we elaborate on the methodologies to extract the numerical responses. There are many different level of information that we can use to characterize numerical responses from LLMs. When estimating empirical distributions using point estimates, it suffices to use greedy decoding, since it represents the best guess that a model would give on average. However, in some cases, one may be interested in more faithful estimates of the empirical probability distribution for each uncertainty expression. To that end, one can take advantage of the information available in different platforms, which we describe next.
C.1 Full Next-Token Probability Information
By definition, our task elicits a numerical response from LLMs, which resembles the setup in verbalized confidence (Tian et al., 2023). The adoption of single digit tokenization (Singh and Strouse, 2024) by autoregressive models (e.g., Gemma (2B), LLama3 (70B), and OLMo (7B)) creates some challenges in the computation of numerical responses non-trivial for autoregressive models. In practice, due to the left-to-right nature of LLMs, single digit tokenization implies that the probability of a number between $[0,9]$ is always greater or equal to the probability of any number in $[10,100]$. To circumvent this problem, we report the corrected probability during our experiments as follows:
$p_{\mathrm{model}}(y_{t}=i|x)-\sum_{j=0}^{9}p_{\mathrm{model}}(y_{t}=i,y_{t+1}=j|x),$
where $x$ is a prompt and $j\in[0,9]$. Intuitively, this means that we are computing the probability of $i\in[0,9]$ and no other number following it. The details of what a number is change with tokenizer implementation. Unlike traditional greedy decoding, we condition the selection of the arg-max prediction to the set of strings representing the numbers between $[0,100]$ (followed by no other number).
C.2 Partial Next-Token Probability Information
In order to work, this method requires two properties to be satisfied: (1) numbers between 0 and 100 were encoded with unique tokens (i.e., there are 101 unique integers that represent each individual token), and (2) exponentiating the log probabilities returned by the black-box API must lead to a valid probability distribution (i.e., numbers obtained for different prompts will be comparable to one another). For OpenAI models, the first requirement is satisfied.
Selecting the greedy prediction: Unlike traditional greedy decoding, we condition the selection of the arg-max prediction over numbers the top-k (k=20 for OpenAI). That is, we select the most likely number that is present in the top-20 predicted tokens.
Estimating conditional distributions using probabilistic decoding: In Section 5.4, we use the information available in the top-20 tokens made available by OpenAI models. Like before we bin the predictions into 21 bins, defined from 0 to 100 in increments of 5. To ensure that we have a valid probability definition, we consider an additional bin (“-1”) that accumulates the probability of not having being a number in the top-20.
(a) Total probability mass assigned to a number across models and settings.
(b) Probability mass of greedy prediction.
Figure 12: Differences in the probability mass as determined by OpenAI models on the top-20 tokens. We report these values across all statements (n=840) in the verifiable and non-verifiable settings. We observe that numbers account for the majority of probability mass in ChatGPT and GPT-4. Upon analysis we found lower probability mass assigned by GPT-4 to be correlated with the appearance of the words “Given”, “The”, and “And”.
In Section 4.2 we briefly mention the use of different methodologies to extract numerical responses from LLMs. The following describes the list of methodologies used for each LLM:
  • Greedy Sampling: Gemini, LLama3 (70B), Mixtral 8x7B, Mixtral 8x22B. We opt for using greedy sampling as opposed to standard sampling due to budget constraints. Given the nature of our experiments, faithfully estimating the empirical distributions over the 100 numbers would require hundreds or thousands of calls. These calls are time-consuming and costly. We believe that using decoding is still representative of how a model would behave in most cases.
  • Full next-token probability distribution: LLama3 (8B), Gemma (2B), OLMo (7B). We found these models to be particularly brittle to the prompts.
  • Next-token probability distribution: ChatGPT, GPT-4, GPT-4o. As of June 2024, OpenAI models only provide access to the next-token probabilities of the top-20 tokens. During the experiments in Section 5.4, we collect the information about the top 20 numbers
Appendix D Additional Results
In this section, we report additional results, including visualization of the empirical distributions for models and humans in Section D.1, additional measurements of similarity between non-verifiable and verifiable distributions in Section D.2, metrics discriminated by uncertainty expression (whenever applicable) in Sections D.3 and D.4.
D.1 Histograms
Figure 13 depicts the empirical distributions for the non-verifiable experiments.
(a) Humans
(b) Random Baseline
(c) Gemma-1.1-2b-it
(d) OLMo-7B-instruct
(e) gemini-pro
(f) GPT-4o
(g) ChatGPT
(h) GPT-4
(i) LLama3 (8B)
(j) LLama3 (70B)
(k) Mixtral 8x7B
(l) Mixtral 8x22B
Figure 13: Empirical distributions of numerical probabilities per uncertainty expression in the non-verifiable setting. For each uncertainty expression (row), the empirical distribution is computed based on n=60 datapoints (for LLMs) and n=188 (for humans). We observe that larger, newer LLMs match the mode of human behavior.
D.2 Summary Metrics
For a more complete understanding of the differences among the distributions, we report distance metrics in Table 4.
  • Proportional Agreement: proposed in Section 4.3, measures the overall agreement between an agent’s and a reference (population) distribution. We use the results of the human studies in the non-verifiable setting as our reference distribution throughout the whole paper.
  • Mean Absolute Error: proposed in Section 4.3, measures the average agreement across uncertainty expressions between a agent’s distribution and a reference (population) distribution. In this case, we also use the human results from the non-verifiable setting as our reference distribution throughout the paper.
  • Wasserstein Distance: computed using scipy.stats.wasserstein_distance, measures the distance between two conditional distributions.
Table 4: Summary metrics averaged across uncertainty expressions. All metrics are computed with respect to the reference distribution in the non-verifiable setting (Human + NV). “PA” reports the general agreement between LLMs and the mode of the human distribution, reported in percentages. “MAE” reports the absolute error between the mean responses of LLMs and those of humans. Wasserstein-1 computes the distance between LLM and the baseline distribution.
Avg PA ($\uparrow$) Avg MAE   ($\downarrow$) Avg Wasserstein-1 ($\downarrow$)
NV V NV V NV V
Human Mode 27.6 27.6
Individual 17.6 16.7 8.91 9.35 12.35 12.99
[…]
Gemma (2B) 8.1 6.6 20.17 24.33 22.13 25.89
Llama3 8B 17.8 10.1 11.99 16.59 14.11 18.35
Llama3 (70B) 23.6 18.8 5.56 13.73 9.94 16.39
Mixtral 8x7B 21.8 15.2 5.88 12.32 8.88 15.93
Mixtral 8x22B 21.8 18.6 7.20 9.78 10.78 12.05
Gemini 25.4 21.3 4.09 7.23 9.24 9.78
ChatGPT 19.7 15.3 6.80 8.57 9.26 12.74
GPT-4 24.4 22.1 4.64 3.84 9.96 6.88
GPT-4o 18.9 15.2 5.58 7.05 10.34 9.96
D.3 Proportional Agreement
Tables 5 and 6 report the proportional agreement (PA) metric discriminated by uncertainty expression in the non-verifiable and verifiable settings, respectively. The results are reported in the filtered pool of human participants.
Table 5: Proportional Agreement (PA) of models in the non-verifiable setting.
Humans Mode Humans Individual GPT-4o GPT-4 ChatGPT Llama3 Mixtral 8x22B Gemini DBRX Random
Avg PA 27.6 17.6 18.9 24.4 19.7 23.6 21.8 25.4 14.2 5.1
[…]
highly likely 34.6 22.5 17.4 22.1 16.1 8.3 23.9 30.0 7.8 6.5
very likely 28.7 17.7 19.1 19.5 13.6 14.4 15.1 21.9 14.8 5.1
probable 16.0 11.1 14.9 14.3 7.6 14.8 15.7 14.1 10.3 5.1
likely 20.2 12.9 16.5 16.5 8.2 19.7 16.5 18.1 8.7 5.0
somewhat likely 20.2 13.5 12.8 17.9 6.8 16.5 18.2 16.2 6.3 5.7
somewhat unlikely 22.3 14.0 18.0 19.3 21.8 22.3 18.4 19.9 15.6 4.9
uncertain 35.1 16.9 35.1 35.1 33.1 35.1 35.1 35.1 35.1 3.4
possible 18.1 10.7 6.5 15.4 14.2 15.0 8.6 15.5 15.6 5.1
unlikely 19.1 13.5 11.1 18.3 15.9 16.8 10.1 16.5 14.1 4.8
not likely 18.1 12.6 11.3 17.1 16.7 17.7 16.1 18.1 10.7 5.7
doubtful 19.7 11.7 14.2 12.9 13.2 17.5 16.4 19.7 11.8 5.8
very unlikely 38.8 23.4 27.8 37.9 26.5 38.8 22.3 36.9 11.5 3.0
highly unlikely 35.1 23.5 24.8 35.1 25.7 35.1 28.7 32.8 10.9 3.8
Table 6: Proportional Agreement (PA) of models in the verifiable setting.
Humans Mode Humans Individual GPT-4o GPT-4 GPT-3.5 Llama3 Mixtral 8x22B Gemini DBRX Random
Avg PA 27.6 16.7 15.2 22.1 15.3 18.8 18.6 21.3 13.6 5.1
[…]
highly likely 34.6 21.5 14.7 22.0 13.0 17.4 24.1 22.8 15.5 6.5
very likely 28.7 17.7 17.2 17.2 10.1 14.9 14.9 17.6 12.7 5.1
probable 16.0 11.4 11.3 11.5 3.8 6.8 5.8 9.9 6.6 5.1
likely 20.2 12.7 12.7 12.4 7.3 10.7 9.2 11.3 6.4 5.0
somewhat likely 20.2 13.3 7.0 12.1 6.5 8.6 9.0 10.0 7.6 5.7
somewhat unlikely 22.3 13.0 15.0 18.8 15.4 17.3 11.8 15.4 14.7 4.9
uncertain 35.1 15.9 30.3 34.2 25.1 34.0 33.4 34.0 31.8 3.4
possible 18.1 9.4 6.1 10.2 5.8 4.0 4.6 6.7 6.9 5.1
unlikely 19.1 13.2 10.6 16.8 13.6 15.0 9.4 14.3 12.6 4.8
not likely 18.1 11.7 10.9 15.0 12.8 15.5 13.3 15.7 12.7 5.7
doubtful 19.7 10.0 11.6 11.5 6.2 15.2 13.8 16.4 11.9 5.8
very unlikely 38.8 21.8 18.8 34.1 24.2 35.9 22.3 36.2 12.8 3.0
highly unlikely 35.1 22.7 17.2 32.8 22.1 33.3 28.7 33.3 14.0 3.8
D.4 Mean Response
Figure 14 illustrates the mean rated probability metric discriminated by uncertainty expressions across the non-verifiable, as well as the true and false verifiable statements.
(a) almost certain
(b) highly likely
(c) highly unlikely
(d) possible
(e) very likely
(f) very unlikely
(g) probable
(h) likely
(i) not likely
(j) unlikely
(k) somewhat likely
(l) somewhat unlikely
(m) uncertain
(n) doubtful
Figure 14: Mean response (and $95\%$ confidence intervals) of verifiable statements discriminated by truthfulness of the statement across all 14 evaluated uncertainty expressions.
Appendix E Generalization results
Diversity of grammatical and semantic structures is an important component of current evaluation practices in LLMs (Selvam et al., 2023; Seshadri et al., 2022), since it helps ensure that obtained results are not an artifact of the evaluation methodology and/or benchmarks used. The experiments described in the main paper were carefully crafted to cover various topics and situations where uncertainty expressions could be used. To further strengthen our analysis and validate our findings, we simultaneously run collect models perceptions of uncertainty expressions using a larger dataset. This dataset by the authors based on the AI2-Arc test set (Clark et al., 2018) — a popular question-answering dataset consisting of genuine grade-school level, multiple-choice science questions. Not only has this dataset been recently used to measure commonsense reasoning of current state-of-the-art LLMs (Jiang et al., 2023; Achiam et al., 2024; Beeching et al., 2023), but it is also composed of easier questions, a key aspect to our verifiable experiment setup.
The creation of this dataset mirrors the procedure described in Section 4. We manually repurposed 200 question-answer pairs from AI2 Arc (100 from the easy set and another 100 from the challenge set). For every statement, the authors produce a true statement and a false statement using the available information about the correct and incorrect multiple choices. The final dataset consists of 200 true statements and 200 false statements.
To determine distributional differences between the conditional distributions obtained in the main paper and the ones obtained in the generalization set, we compare the Wasserstein-1 distance of the two empirical distributions. These values are reported in Table 7. In general, we find models that performed worse in the main paper, including Gemma (2B) and OLMo (7B), to exhibit the largest distributional differences with Wasserstein distances of 48.5 and 13.1 when averaged over uncertainty expressions. ChatGPT, LLama3 (70B), and Mixtral models all exhibit higher differences in expressions of higher certainty, e.g., “highly likely”, “probable”, “possible”. On the other hand, the two GPT-4 models, as well as Gemini (Pro) suffer the least changes distributionally (1.9, 1.8, and 4.2 Wasserstein-1 distances on average, respectively), suggesting that these models were robust to changes in the statements.
Table 7: Difference between the empirical distributions of each model in the original and the generalization datasets. Differences are computed for each uncertainty expression in terms of the Wasserstein distance between the two distributions. Smaller values imply similar distributions. We observe minimal changes in models like GPT-4o, GPT-4, and Gemini and larger differences in models like OLMo (7B), Gemma (2B), and ChatGPT.
GPT-4o GPT-4 ChatGPT Gemini Llama3 (70B) Mixtral 8x7B Mixtral 8x22B OLMo (7B) Gemma (2B)
Avg 1.9 1.8 9.1 4.2 8.8 7.6 3.7 13.1 48.5
[…]
highly likely 1.5 1.3 12.8 5.1 17.1 11.2 2.2 18.9 55.1
very likely 1.6 1.8 10.6 7.0 16.0 13.3 1.8 18.4 53.6
likely 4.3 2.9 10.9 8.1 13.8 10.4 9.0 14.3 51.4
probable 3.2 3.2 17.2 5.8 14.4 16.0 10.6 12.8 50.7
somewhat likely 2.6 4.5 5.1 4.4 8.5 7.8 11.1 8.0 42.2
possible 4.8 3.4 14.6 7.5 11.6 11.2 6.6 15.0 48.6
uncertain 0.5 0.7 12.8 1.5 1.6 2.2 1.7 10.9 47.9
somewhat unlikely 2.3 0.4 3.6 2.0 2.1 4.2 1.6 14.0 34.2
unlikely 0.7 1.1 4.6 3.4 4.3 6.3 1.3 8.4 42.3
not likely 0.7 1.8 8.3 2.4 4.7 0.6 1.5 13.6 48.7
doubtful 0.2 1.5 14.6 2.5 5.1 6.6 1.9 24.7 47.1
very unlikely 1.2 0.8 2.1 1.4 2.6 1.1 0.2 3.1 50.7
highly unlikely 1.0 1.0 3.0 2.1 3.3 1.8 0.2 7.1 54.4
In the main paper, we find it surprising that LLMs perception abilities differ significantly based on whether the uncertainty expressions are referring to someone’s belief in a true or false statement. To test the generalization of this finding in a larger (and different) dataset, we repeat the same analysis and compare the observed mean response differences with that of humans obtained in the original setting (see Figures 15 and 16). We observe that in absolute sense the differences are smaller than those observed in the original setting, but that models are affected by this knowledge gap to a greater extent than humans.
Figure 15: Mean numerical response (and $95\%$ confidence intervals) for verifiable statements across true and false statements averaged over the uncertainty expressions in the generalization set. Similarly to the main experiment, we observe large absolute differences between the mean numerical responses produced by LLMs in true statements and false statements. However, the magnitude of these changes is slightly lower than in the original experiment.
(a) almost certain
(b) highly likely
(c) highly unlikely
(d) possible
(e) very likely
(f) very unlikely
(g) probable
(h) likely
(i) not likely
(j) unlikely
(k) somewhat likely
(l) somewhat unlikely
(m) uncertain
(n) doubtful
Figure 16: Mean response (and $95\%$ confidence intervals) of verifiable statements across true and false statements for the 14 evaluated uncertainty expressions in the generalization set.
Appendix F Probabilistic decoding
Table 8 summarizes the different metrics in the non-verifiable setting, when using probabilistic decoding with OpenAI models, i.e., temperature = 1. The probability distributions are obtained using the methodology described in Appendix C.
Table 8: Summary metrics average across uncertainty expressions using probabilistic decoding (temperature=1) in both non-verifiable and verifiable settings.
Avg PA ($\uparrow$) Avg MAE   ($\downarrow$) Avg Wasserstein-1 ($\downarrow$)
NV V NV V NV V
Human Mode 27.6 27.6
Individual 17.6 16.7 8.91 9.35 12.35 12.99
[…]
GPT4 24.4 21.4 4.62 4.00 9.78 6.72
GPT4o 12.9 8.7 19.01 26.07 19.69 26.14
(a) almost certain
(b) highly likely
(c) highly unlikely
(d) possible
(e) very likely
(f) very unlikely
(g) probable
(h) likely
(i) not likely
(j) unlikely
(k) somewhat likely
(l) somewhat unlikely
(m) uncertain
(n) doubtful
Figure 17: Mean response (and $95\%$ confidence intervals) of verifiable statements across true and false statements for the 14 evaluated uncertainty expressions, when using probabilistic decoding (i.e., temperature=1).